# Is Epistemic Uncertainty Faithfully Represented by Evidential Deep Learning Methods?

### v2

@@ 4.2 Results
Figure 2 shows the estimated reference distribution and the fitted second-order models for different sample sizes. For outer loss minimization without regularization, the conclusions of Theorem 3.3 are nicely reflected in the results: A distribution very close to a Dirac is obtained, with very narrow confidence bounds around the mean prediction. When regularization is added, the confidence bounds become wider, but still deviate substantially from those obtained for the reference distribution, especially in the part of the input space where no data has been observed. Hence the epistemic uncertainty quantification by the second-order models is arbitrary, and this phenomenon occurs over various sample sizes. For inner loss minimization, a higher epistemic uncertainty is estimated for the example run that is shown, but here, too, the estimated second-order distributions deviate substantially from the reference. However, in accordance with Theorem 3.2, we would like to emphasize that the parameters of the second-order distribution are unidentifiable for inner loss minimization. For a different run of the optimization algorithm on the same training data, we observed that the results may look very different.
To analyze some of these observations more deeply, we perform a second experiment in which we train the models with more epochs (up to 150 000) and 40 runs on the same training dataset. The variation between different runs comes from the uniform initialization of the network parameters. In Figure 3 the estimated second-order parameters $\hat{\alpha}$ and $\hat{\beta}$ as a function of the number of epochs for different runs are displayed. For outer loss minimization, one can see that no convergence occurred after 15 000 epochs. The estimates of $\alpha$ and $\beta$, averaged over the instance space, keep growing in magnitude, which is in accordance with Theorem 3.3. When regularization is added, $\alpha$ and $\beta$ grow in a much smaller rate, thereby also solving the convergence issues, as expected due to Theorem 3.4. For inner loss minimization, one can observe large standard deviations in the estimates for $\alpha$ and $\beta$ when the Adam optimizer is running multiple times on the same training dataset. This is expected because of Theorem 3.2. Adding some regularization substantially decreases this variation.
@@ 5 Conclusion
In this paper, we proposed the comparison with a reference distribution as an objective procedure to assess whether evidential deep learning methods represent epistemic uncertainty in a faithful manner. Using this procedure, we showed that epistemic uncertainty is in general not faithfully represented. More specifically, the regularizers that characterize evidential deep learning methods are mainly used to stabilize the corresponding optimization problems, but the resulting measures of epistemic uncertainty cannot be interpreted in a quantitative manner. In many downstream tasks, such as out-of-distribution detection, only relative interpretations of epistemic uncertainty are needed. This probably explains the strong performance of evidential deep learning methods on such tasks. We hypothesize that the OOD scores reported by evidential deep learning methods could be interpreted as density estimates in feature space. This observation was made by Grathwohl et al. (2020); Charpentier et al. (2022), but we believe that more research is needed to draw final conclusions of that kind.
Impact and ethics statement
Machine learning systems are increasingly prevalent in high-stake application domains with a direct impract on humans, such as medical diagnostics, legal decisions, self-driving cars, fraud detection, natural disaster forecasting, etc. The profound consequences of inaccurate predictions in these areas emphasize the necessity of assessing the certainty of machine-learned forecasts. The current upswing in research on uncertainty quantification in the machine learning community underscores the recognition of this crucial aspect.
@@ Appendix A Basic properties of exponential family members
This is the probability mass function of the Categorical distribution with parameters $\vartheta_{1},\vartheta_{2},...,\vartheta_{k}$. In this case $\mathcal{H}_{J}$ does not contain injective $\mathcal{M}\rightarrow\mathbb{P}(\mathcal{Y})$ operators, and $\mathcal{C}_{1}=\mathcal{C}_{J}$ because the two hypothesis spaces contain Categorical distributions. When normal-inverse Gamma distributions are chosen for $\mathcal{H}_{2}$, the predictive distribution is derived by Amini et al. [2020]:
@@ Appendix G Additional experiments
For the regression experiments we use the same setup as introduced in Amini et al. [2020], and critically analyzed by Meinert and Lavin [2022]. That is, we generate datasets $\{(x_{i},x_{i}^{3}+\epsilon)\}_{i=1}^{N}$ of different sample sizes, where the instances $x_{i}\in U([-4,4])$ are uniformly distributed and $\epsilon\sim\mathcal{N}(0,\sigma^{2}=9)$. Even though the setup is homoscedastic, we use negative log-likelihood loss for the first-order learner, while assuming an underlying normal distribution, making $\sigma$ as well as $\mu$ learnable parameters. This choice is made because the second-order risk minimization methods also assume heteroskedastic noise, so in this way the comparison is more fair.
@@
Figure 5 shows the learned parameters as a function of the number of epochs (up to 10000) for 40 runs with random initialization of the neural network. For the unregularized case we see that the network learns to increase $\nu$ over time, thereby decreasing the variance of $\widehat{\mu}$. For the inner expectation minimization, we obtain similar results as in Amini et al. [2020]. Note here that for comparability to the first experiment we use negative entropy instead of the proposed evidence based regularizor. For the unregularized case, the distribution of $\widehat{\sigma^{2}}$, whose expectation is an indicator of the aleatoric uncertainty, is more peaked in the center of region where the model sees data, which is in accordance to the results of Meinert and Lavin [2022]. Again the issue of non-identifiability of the learned parameters arises, and we obtained a high variability in the results for different runs.

### v3

@@ 4.2 Results
Figure 2 shows the mean prediction as well as the $95\%$ confidence levels for the reference model and the fitted second-order models, trained using different sample sizes. For outer loss minimization without entropy regularization, the conclusions of Theorem 3.3 are nicely reflected in the results: The narrow confidence bounds around the mean prediction indicate a distribution very close to a Dirac delta. When regularization is added, the confidence bounds become wider, but still deviate substantially from those obtained for the reference distribution, especially in the part of the input space where no data has been observed. Hence the epistemic uncertainty quantification by the second-order models is arbitrary and tightly bound to the choice of the regularization strength $\lambda$, and this phenomenon occurs over various sample sizes. For inner loss minimization, a higher epistemic uncertainty is estimated for the example run that is shown, but here, too, the estimated second-order distributions deviate substantially from the reference. However, in accordance with Theorem 3.2, we would like to emphasize that the parameters of the second-order distribution are unidentifiable for inner loss minimization. For a different run of the optimization algorithm on the same training data, we observed that the results may look very different.
In order to analytically evaluate the distance between the reference distribution and the second-order distributions, the empirical Wasserstein-$1$ distance was calculated per instance and for different $\lambda$ values. Figure 3 shows the results: Both in regions with and without training data, a relatively high distance to the reference distribution is observed. Using different orders of magnitude for $\lambda$ does not seem to significantly change that pattern.
To analyze some of these observations more deeply, we perform a second experiment in which we train the models with more epochs (up to 15 000) and 40 runs on the same training dataset. The variation between different runs comes from the uniform initialization of the network parameters. In Figure 4 the estimated second-order parameters $\hat{\alpha}$ and $\hat{\beta}$ as a function of the number of epochs for different runs are displayed. For outer loss minimization, one can see that no convergence occurred after 15 000 epochs. The estimates of $\alpha$ and $\beta$, averaged over the instance space, keep growing in magnitude, which is in accordance with Theorem 3.3. When regularization is added, $\alpha$ and $\beta$ grow in a much smaller rate, thereby also solving the convergence issues, as expected due to Theorem 3.4. For inner loss minimization, one can observe large standard deviations in the estimates for $\alpha$ and $\beta$ when the Adam optimizer is running multiple times on the same training dataset. This is expected because of Theorem 3.2. Adding some regularization substantially decreases this variation.
@@ 5 Conclusion
In this paper, we proposed the comparison with a reference distribution as an objective procedure to assess whether evidential deep learning methods represent epistemic uncertainty in a faithful manner. Using this procedure, we showed that epistemic uncertainty is in general not faithfully represented. More specifically, the regularizers that characterize evidential deep learning methods are mainly needed to stabilize the corresponding optimization problems, thereby yielding an ”uncertainty budget” that cannot be exceeded. Conversely, in first-order risk minimization, regularization is mainly used to prevent overfitting by controlling the bias-variance trade-off. The resulting measures of epistemic uncertainty cannot be interpreted in a quantitative manner. In many downstream tasks, such as out-of-distribution detection, only relative interpretations of epistemic uncertainty are needed. This probably explains the strong performance of evidential deep learning methods on such tasks. We hypothesize that the OOD scores reported by evidential deep learning methods could be interpreted as density estimates in feature space. This observation was made by Grathwohl et al. (2020); Charpentier et al. (2022), but we believe that more research is needed to draw final conclusions of that kind.
Impact statement
Machine learning systems are increasingly prevalent in high-stake application domains with a direct impact on humans, such as medical diagnostics, legal decisions, self-driving cars, fraud detection, natural disaster forecasting, etc. The profound consequences of inaccurate predictions in these areas emphasize the necessity of assessing the certainty of machine-learned forecasts. The current upswing in research on uncertainty quantification in the machine learning community underscores the recognition of this crucial aspect.
@@
Acknowledgements
M.J. and W.W. received funding from the Flemish Government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme.
@@ Appendix A Basic properties of exponential family members
This is the probability mass function of the Categorical distribution with parameters $\vartheta_{1},\vartheta_{2},...,\vartheta_{k}$. In this case $\mathcal{H}_{J}$ does not contain injective $\mathcal{M}\rightarrow\mathbb{P}(\mathcal{Y})$ operators, and $\mathcal{C}_{1}=\mathcal{C}_{J}$ because the two hypothesis spaces contain Categorical distributions.
When normal-inverse Gamma distributions are chosen for $\mathcal{H}_{2}$, the predictive distribution is derived by Amini et al. [2020]:
@@ Appendix G Additional experiments
For the regression experiments we use the same setup as introduced in [Amini et al., 2020], and critically analyzed by Meinert and Lavin [2022]. That is, we generate datasets $\{(x_{i},x_{i}^{3}+\epsilon)\}_{i=1}^{N}$ of different sample sizes, where the instances $x_{i}\in U([-4,4])$ are uniformly distributed and $\epsilon\sim\mathcal{N}(0,\sigma^{2}=9)$. Even though the setup is homoscedastic, we use negative log-likelihood loss for the first-order learner, while assuming an underlying normal distribution, making $\sigma$ as well as $\mu$ learnable parameters. This choice is made because the second-order risk minimization methods also assume heteroskedastic noise, so in this way the comparison is more fair.
@@
Figure 6 shows the learned parameters as a function of the number of epochs (up to 10000) for 40 runs with random initialization of the neural network. For the unregularized case we see that the network learns to increase $\nu$ over time, thereby decreasing the variance of $\widehat{\mu}$. For the inner expectation minimization, we obtain similar results as in [Amini et al., 2020]. Note here that for comparability to the first experiment we use negative entropy instead of the proposed evidence based regularizor. For the unregularized case, the distribution of $\widehat{\sigma^{2}}$, whose expectation is an indicator of the aleatoric uncertainty, is more peaked in the center of region where the model sees data, which is in accordance to the results of [Meinert and Lavin, 2022]. Again the issue of non-identifiability of the learned parameters arises, and we obtained a high variability in the results for different runs.